### v2

@@ 4 A simple example: ${\sf N}=5$
Having explained our algorithm in §3 in full generality, we now return to the set-up described in §1 and §2 to exemplify the application of the algorithm to the SAC. Specifically, we will use it to compute the subset of $\mathscr{R}_{{}_{\text{KC}}}^{5}$ of “genuine $5$-party” ERs, i.e., those that are not lifts of element of $\mathscr{R}_{{}_{\text{KC}}}^{{\sf N}}$ for some ${\sf N}<5$. These ERs are well-known,464646 It was shown in He:2022bmi that $\mathscr{R}_{{}_{\text{KC}}}^{5}=\mathscr{R}_{{}_{\text{SSA}}}^{5}$, and $\mathscr{R}_{{}_{\text{SSA}}}^{5}$ was computed in Hernandez-Cuenca:2019jpv . and our goal here is simply to demonstrate how our algorithm works in a particular example. In the next section, we will use the algorithm to compute the much more involved $\mathscr{R}_{{}_{\text{KC}}}^{6}$ (with the main goal of finding $\mathscr{R}_{{}_{\text{SSA}}}^{6}$), and there we will only present the result.
@@
Moreover, we will use two of the variations mentioned in §3.4 when implementing the basic algorithm presented in §3.3. The first variation is that we will use the closure operator $\text{cl}_{\text{LD}}$ instead of $\text{cl}_{\text{FD}}$ in the main subroutine. Each triplet will then correspond to a KC-subspace474747 As the reader might guess, a KC-subspace is precisely the D-subspace defined in §3 when the poset is the MI-poset. Similarly, KC-faces and KC-ERs are respectively the D-faces and D-ERs defined above when the poset is the MI-poset. rather than a KC-face. The reason for this choice is that computing $\text{cl}_{\text{LD}}$ is typically much faster than computing $\text{cl}_{\text{FD}}$, and empirical observation suggests that most instances of KC-subspaces are in fact spanned by KC-faces. Thus, using $\text{cl}_{\text{FD}}$ instead of $\text{cl}_{\text{LD}}$ would not give a significant advantage, while dramatically slowing down the computation. We will continue to use the same strategy in the next section.
@@ 5 Results for ${\sf N}=6$
Having exemplified the workings of our algorithm for the ${\sf N}=5$ case, we now present the results of the algorithm applied to finding the KC-ERs of SAC6, primarily focusing on the ones that satisfy SSA.484848 Further details are presented in notes , in particular their adjacency relation, their characterization in terms of the $\beta$-sets of Hubeny:2024wip , and our procedure to obtain graph models that realize most of them. The algorithm produced 220 orbits of genuine $6$-party ERs in $\mathscr{R}_{{}_{\text{KC}}}^{6}$, and upon checking SSA, we find that 12 of these in fact violate at least one instance of SSA (cf. §5.1).494949 Unlike the ${\sf N}\leq 5$ computations, we find that for ${\sf N}=6$ there in fact even exist 1-dimensional KC-subspaces which are not extreme rays of SAC6 because they violate at least one instance of SA. An example of such a subspace is that generated by $\vec{{\sf S}}=${4, 2, 3, 2, 2, 3; 6, 7, 6, 6, 7, 5, 4, 4, 5, 5, 5, 6, 4, 5, 5; 7, 8, 8, 9, 9, 9, 8, 8, 9, 9, 7, 7, 8, 6, 7, 5, 7, 8, 8, 7; 9, 9, 8, 10, 7, 9, 9, 8, 6, 9, 9, 10, 8, 11, 10; 7, 6, 6, 7, 6, 8; 4}.
@@
In Table 4, we present the entropy vectors of a representative instance from every orbit in $\mathscr{R}_{{}_{\text{SSA}}}^{6}$ (this data can also be found in the supplementary materials), where we did not include the lifts of the elements of $\mathscr{R}_{{}_{\text{SSA}}}^{6}$ for ${\sf N}<6$. Following the usual convention, the components of each vector are arranged by the cardinality of the subsystems $\mathscr{J}$ (separated by semicolon for ease of identification), and within each group they are arranged in lexicographic order. For convenience, we also color-code each ER by its type: Warm colors (pink, red, orange, yellow) represent holographic ERs (those definitely in $\mathscr{R}_{{}_{\text{Hol}}}$), cold colors (blue, violet) represent ERs that violate at least one HEI (those definitely in $\mathscr{R}_{{}_{\text{SSA}}}\setminus\mathscr{R}_{{}_{\text{Hol}}}$), and green represents the remaining mystery ERs. This color scheme is further subdivided as follows. Pink ERs are realized by star graphs, red ones by other simple505050 By simple tree graphs, we mean graphs which have no cycles and have one boundary vertex per party. tree graphs, and orange and yellow ones by more complicated graphs. The orange ones were known previously hecdata , while the yellow ones are new, illustrating the power of our construction. Notice that since all ERs of the SAC which can be realized by graph models are simultaneously ERs of the HEC, the yellow ERs yield previously unknown ERs of the ${\sf N}=6$ HEC. Finally, among the ERs that violate some HEI, we distinguish the ones which violate MMI by blue and the ones which only violate some higher ${\sf N}=5,6$ HEI by purple.
@@ 5.2 Extreme rays that violate holographic entropy inequalities
As we mentioned above, 52 of the 208 orbits of ERs in $\mathscr{R}_{{}_{\text{SSA}}}^{6}$ violate some known HEI, and it is natural to further categorize these inequalities into three families: monogamy of mutual information (MMI) Hayden:2011ag (including its lifts), ${\sf N}=5$ HEIs Bao:2015bfa ; Cuenca:2019uzx , and ${\sf N}=6$ HEIs Hernandez-Cuenca:2023iqh (the latter list is incomplete, so for this case the number of violations is merely a lower bound). Regarding the instances of MMI, up to permutations and purifications there are only three distinct possibilities for the cardinalities of the arguments of the tripartite information515151 The tripartite information is defined as ${\sf I}_{3}(\underline{\mathscr{I}}:\underline{\mathscr{J}}:\underline{\mathscr{K}})={\sf S}_{\underline{\mathscr{I}}}+{\sf S}_{\underline{\mathscr{J}}}+{\sf S}_{\underline{\mathscr{K}}}-{\sf S}_{\underline{\mathscr{I}}\underline{\mathscr{J}}}-{\sf S}_{\underline{\mathscr{I}}\underline{\mathscr{K}}}-{\sf S}_{\underline{\mathscr{J}}\underline{\mathscr{K}}}+{\sf S}_{\underline{\mathscr{I}}\underline{\mathscr{J}}\underline{\mathscr{K}}}$, and MMI is the statement that ${\sf I}_{3}(\underline{\mathscr{I}}:\underline{\mathscr{J}}:\underline{\mathscr{K}})\leq 0$: (1:1:1), (1:1:2), and (1:2:2). As it turns out, only the last type can be violated.
@@ 5.3 Extreme rays realizable by holographic graph models
Finally, we remark that some of the ERs with more complicated graphs could potentially have an even simpler graph realization. However, Hubeny:2024wip presents a necessary condition for the realizability of an arbitrary entropy vector by a simple tree graph model. We checked this condition for all 220 orbits in $\mathscr{R}_{{}_{\text{KC}}}^{6}$ (including the ones that violate SSA), and the only orbits that satisfy this condition are precisely #1 to #44, which are the ones we realized by simple trees. This is perhaps an indication that the condition found in Hubeny:2024wip might in fact be sufficient as well.
@@ 6 Discussion
Finally, since these questions are specific to faces of the SAC, it would be worthwhile to explore whether the algorithm could be reformulated using the $\beta$-sets technology introduced in Hubeny:2024wip . This formulation takes advantage of additional structural properties of SSA-compatible faces and, if incorporated into the algorithm, might help restrict the search space and improve efficiency.
@@
The weak form of the conjecture proposed in Hernandez-Cuenca:2022pst states that for any given ${\sf N}$, all extreme rays of the ${\sf N}$-party holographic entropy cone can be obtained from subsystem coarse-grainings of extreme rays in $\mathscr{R}_{{}_{\text{Hol}}}^{{\sf N}^{\prime}}$ for some ${\sf N}^{\prime}\geq{\sf N}$. A stronger form of this conjecture asserts that every extreme ray of the HEC can be realized by a graph model with tree topology (although not necessarily simple; that is, not all boundary vertices are required to be labeled by distinct parties).545454 We refer the reader to Hernandez-Cuenca:2022pst for the proof that the strong form implies the weak one. One of the motivations of this work was to generate data that could be used to corroborate these conjectures.
@@ Acknowledgements.
We would like to especially thank David Simmons-Duffin for generously providing us supercomputing time in the early iterations of the algorithm. We would also like to thank Sergio Hernández-Cuenca and Sridip Pal for useful discussions. T.H. is supported by the Heising-Simons Foundation “Observational Signatures of Quantum Gravity” collaboration grant 2021 -2817, the U.S. Department of Energy, Office of Science, Office of High Energy Physics, under Award No. DE-SC0011632, and the Walter Burke Institute for Theoretical Physics. V.H. has been supported in part by the U.S. Department of Energy grant DE-SC0009999 and by funds from the University of California. M.R. acknowledges support from UK Research and Innovation (UKRI) under the UK government’s Horizon Europe guarantee (EP/Y00468X/1), and by funds from the University of California during the early stages of this work. V.H. acknowledges the hospitality of the Kavli Institute for Theoretical Physics (KITP), the Aspen Center for Physics, and the Tsinghua Southeast Asia Center, where part of this work was done. V.H. and M.R. would like to thank the Yukawa Institute for Theoretical Physics for hospitality during the program “Quantum Information, Quantum Matter and Quantum Gravity” during the early stages of this work. All data is available within the paper, as an ancillary file in the arXiv, and on GitHub github . The code used to generate and analyze this data is a combination of a Wolfram Mathematica implementation of the algorithm outlined in this work, and the freely available software Normaliz Normaliz:301 .
@@ Appendix A Comments on the relation between closure operators and symmetries
Notice that in the expression above, the left-hand side is a closed set, while in general the right-hand side need not be if we do not impose any restriction on the choice of cl and $g$555555 For example, in Figure 2 if $g=(23)$ (in cycles notation) and $X=\{1,2\}$, then $\text{cl}_{\text{F}}(g(\{1,2\}))=\text{cl}_{\text{F}}(\{1,3\})=\{1,2,3,4\}\neq g(\text{cl}_{\text{F}}(\{1,2\}))=g(\{1,2\})=\{1,3\}.$ It follows that a basic necessary condition for (60) to hold is that $g$ must be chosen such that the set $\mathcal{L}_{\text{cl}}$ is closed with respect to $g$, i.e., $g$ maps closed sets to closed sets. We can then consider the largest subgroup $G$ of $\text{Sym}(|\mathcal{E}|)$ such that each $g\in G$ has this property.

### v3

@@ 4 A simple example: ${\sf N}=5$
Having explained our algorithm in §3 in full generality, we now return to the set-up described in §1 and §2 to exemplify the application of the algorithm to the SAC. Specifically, we will use it to compute the subset of $\mathscr{R}_{{}_{\text{KC}}}^{5}$ of “genuine $5$-party” ERs, i.e., those that are not lifts of element of $\mathscr{R}_{{}_{\text{KC}}}^{{\sf N}}$ for some ${\sf N}<5$. These ERs are well-known,454545 It was shown in He:2022bmi that $\mathscr{R}_{{}_{\text{KC}}}^{5}=\mathscr{R}_{{}_{\text{SSA}}}^{5}$, and $\mathscr{R}_{{}_{\text{SSA}}}^{5}$ was computed in Hernandez-Cuenca:2019jpv . and our goal here is simply to demonstrate how our algorithm works in a particular example. In the next section, we will use the algorithm to compute the much more involved $\mathscr{R}_{{}_{\text{KC}}}^{6}$ (with the main goal of finding $\mathscr{R}_{{}_{\text{SSA}}}^{6}$), and there we will only present the result.
@@
Moreover, we will use two of the variations mentioned in §3.4 when implementing the basic algorithm presented in §3.3. The first variation is that we will use the closure operator $\text{cl}_{\text{LD}}$ instead of $\text{cl}_{\text{FD}}$ in the main subroutine. Each triplet will then correspond to a KC-subspace464646 As the reader might guess, a KC-subspace is precisely the D-subspace defined in §3 when the poset is the MI-poset. Similarly, KC-faces and KC-ERs are respectively the D-faces and D-ERs defined above when the poset is the MI-poset. rather than a KC-face. The reason for this choice is that computing $\text{cl}_{\text{LD}}$ is typically much faster than computing $\text{cl}_{\text{FD}}$, and empirical observation suggests that most instances of KC-subspaces are in fact spanned by KC-faces. Thus, using $\text{cl}_{\text{FD}}$ instead of $\text{cl}_{\text{LD}}$ would not give a significant advantage, while dramatically slowing down the computation. We will continue to use the same strategy in the next section.
@@ 5 Results for ${\sf N}=6$
Having exemplified the workings of our algorithm for the ${\sf N}=5$ case, we now present the results of the algorithm applied to finding the KC-ERs of SAC6, primarily focusing on the ones that satisfy SSA.474747 Further details are presented in rays-notes , in particular their adjacency relation and their characterization in terms of the $\beta$-sets of beta-sets . Our heuristic procedure to obtain graph model realizations for most of them is described in graph-constructions . The algorithm produced 220 orbits of genuine $6$-party ERs in $\mathscr{R}_{{}_{\text{KC}}}^{6}$, and upon checking SSA, we find that 12 of these in fact violate at least one instance of SSA (cf. §5.1).484848 Unlike the ${\sf N}\leq 5$ computations, we find that for ${\sf N}=6$ there in fact even exist 1-dimensional KC-subspaces which are not extreme rays of SAC6 because they violate at least one instance of SA. An example of such a subspace is that generated by $\vec{{\sf S}}=${4, 2, 3, 2, 2, 3; 6, 7, 6, 6, 7, 5, 4, 4, 5, 5, 5, 6, 4, 5, 5; 7, 8, 8, 9, 9, 9, 8, 8, 9, 9, 7, 7, 8, 6, 7, 5, 7, 8, 8, 7; 9, 9, 8, 10, 7, 9, 9, 8, 6, 9, 9, 10, 8, 11, 10; 7, 6, 6, 7, 6, 8; 4}.
@@
In Table 4, we present the entropy vectors of a representative instance from every orbit in $\mathscr{R}_{{}_{\text{SSA}}}^{6}$ (this data can also be found in the supplementary materials), where we did not include the lifts of the elements of $\mathscr{R}_{{}_{\text{SSA}}}^{6}$ for ${\sf N}<6$. Following the usual convention, the components of each vector are arranged by the cardinality of the subsystems $\mathscr{J}$ (separated by semicolon for ease of identification), and within each group they are arranged in lexicographic order. For convenience, we also color-code each ER by its type: Warm colors (pink, red, orange, yellow) represent holographic ERs (those definitely in $\mathscr{R}_{{}_{\text{Hol}}}$), cold colors (blue, violet) represent ERs that violate at least one HEI (those definitely in $\mathscr{R}_{{}_{\text{SSA}}}\setminus\mathscr{R}_{{}_{\text{Hol}}}$), and green represents the remaining mystery ERs. This color scheme is further subdivided as follows. Pink ERs are realized by star graphs, red ones by other simple494949 By simple tree graphs, we mean graphs which have no cycles and have one boundary vertex per party. tree graphs, and orange and yellow ones by more complicated graphs. The orange ones were known previously hecdata , while the yellow ones are new, illustrating the power of our construction. Notice that since all ERs of the SAC which can be realized by graph models are simultaneously ERs of the HEC, the yellow ERs yield previously unknown ERs of the ${\sf N}=6$ HEC. Finally, among the ERs that violate some HEI, we distinguish the ones which violate MMI by blue and the ones which only violate some higher ${\sf N}=5,6$ HEI by purple.
@@ 5.2 Extreme rays that violate holographic entropy inequalities
As we mentioned above, 52 of the 208 orbits of ERs in $\mathscr{R}_{{}_{\text{SSA}}}^{6}$ violate some known HEI, and it is natural to further categorize these inequalities into three families: monogamy of mutual information (MMI) Hayden:2011ag (including its lifts), ${\sf N}=5$ HEIs Bao:2015bfa ; Cuenca:2019uzx , and ${\sf N}=6$ HEIs Hernandez-Cuenca:2023iqh (the latter list is incomplete, so for this case the number of violations is merely a lower bound). Regarding the instances of MMI, up to permutations and purifications there are only three distinct possibilities for the cardinalities of the arguments of the tripartite information505050 The tripartite information is defined as ${\sf I}_{3}(\underline{\mathscr{I}}:\underline{\mathscr{J}}:\underline{\mathscr{K}})={\sf S}_{\underline{\mathscr{I}}}+{\sf S}_{\underline{\mathscr{J}}}+{\sf S}_{\underline{\mathscr{K}}}-{\sf S}_{\underline{\mathscr{I}}\underline{\mathscr{J}}}-{\sf S}_{\underline{\mathscr{I}}\underline{\mathscr{K}}}-{\sf S}_{\underline{\mathscr{J}}\underline{\mathscr{K}}}+{\sf S}_{\underline{\mathscr{I}}\underline{\mathscr{J}}\underline{\mathscr{K}}}$, and MMI is the statement that ${\sf I}_{3}(\underline{\mathscr{I}}:\underline{\mathscr{J}}:\underline{\mathscr{K}})\leq 0$: (1:1:1), (1:1:2), and (1:2:2). As it turns out, only the last type can be violated.
@@ 5.3 Extreme rays realizable by holographic graph models
Finally, we remark that some of the ERs with more complicated graphs could potentially have an even simpler graph realization. However, beta-sets presents a necessary condition for the realizability of an arbitrary entropy vector by a simple tree graph model. We checked this condition for all 220 orbits in $\mathscr{R}_{{}_{\text{KC}}}^{6}$ (including the ones that violate SSA), and the only orbits that satisfy this condition are precisely #1 to #44, which are the ones we realized by simple trees. This is perhaps an indication that the condition found in beta-sets might in fact be sufficient as well.
@@ 6 Discussion
Finally, since these questions are specific to faces of the SAC, it would be worthwhile to explore whether the algorithm could be reformulated using the $\beta$-sets technology introduced in beta-sets . This formulation takes advantage of additional structural properties of SSA-compatible faces and, if incorporated into the algorithm, might help restrict the search space and improve efficiency.
@@
The weak form of the conjecture proposed in Hernandez-Cuenca:2022pst states that for any given ${\sf N}$, all extreme rays of the ${\sf N}$-party holographic entropy cone can be obtained from subsystem coarse-grainings of extreme rays in $\mathscr{R}_{{}_{\text{Hol}}}^{{\sf N}^{\prime}}$ for some ${\sf N}^{\prime}\geq{\sf N}$. A stronger form of this conjecture asserts that every extreme ray of the HEC can be realized by a graph model with tree topology (although not necessarily simple; that is, not all boundary vertices are required to be labeled by distinct parties).525252 We refer the reader to Hernandez-Cuenca:2022pst for the proof that the strong form implies the weak one. One of the motivations of this work was to generate data that could be used to corroborate these conjectures.
@@ Acknowledgements.
We would like to especially thank David Simmons-Duffin for generously providing us supercomputing time in the early iterations of the algorithm. We would also like to thank Sergio Hernández-Cuenca and Sridip Pal for useful discussions. T.H. is supported by the Heising-Simons Foundation “Observational Signatures of Quantum Gravity” collaboration grant 2021 -2817, the U.S. Department of Energy, Office of Science, Office of High Energy Physics, under Award No. DE-SC0011632, and the Walter Burke Institute for Theoretical Physics. V.H. has been supported in part by the U.S. Department of Energy grant DE-SC0009999 and by funds from the University of California. M.R. acknowledges support from UK Research and Innovation (UKRI) under the UK government’s Horizon Europe guarantee (EP/Y00468X/1), and by funds from the University of California during the early stages of this work. V.H. acknowledges the hospitality of the Kavli Institute for Theoretical Physics (KITP), the Aspen Center for Physics, and the Tsinghua Southeast Asia Center, where part of this work was done. V.H. and M.R. would like to thank the Yukawa Institute for Theoretical Physics for hospitality during the program “Quantum Information, Quantum Matter and Quantum Gravity” during the early stages of this work. All data is available within the paper, as an ancillary file in the arXiv, and on GitHub github . The code used to generate and analyze this data is a combination of a Wolfram Mathematica implementation of the algorithm outlined in this work, and the freely available software Normaliz Normaliz:301 . For the purpose of open access, the authors have applied a Creative Commons Attribution (CC BY) licence to any Author Accepted Manuscript version arising from this submission.
@@ Appendix A Comments on the relation between closure operators and symmetries
Notice that in the expression above, the left-hand side is a closed set, while in general the right-hand side need not be if we do not impose any restriction on the choice of cl and $g$535353 For example, in Figure 2 if $g=(23)$ (in cycles notation) and $X=\{1,2\}$, then $\text{cl}_{\text{F}}(g(\{1,2\}))=\text{cl}_{\text{F}}(\{1,3\})=\{1,2,3,4\}\neq g(\text{cl}_{\text{F}}(\{1,2\}))=g(\{1,2\})=\{1,3\}.$ It follows that a basic necessary condition for (60) to hold is that $g$ must be chosen such that the set $\mathcal{L}_{\text{cl}}$ is closed with respect to $g$, i.e., $g$ maps closed sets to closed sets. We can then consider the largest subgroup $G$ of $\text{Sym}(|\mathcal{E}|)$ such that each $g\in G$ has this property.